-4x-15=57
5x+2y=11
How do you solve systems equations?

Answer by Alan3354(69443)

-4x-15=57
-4x = 72
x = -18
--------
5x+2y=11
5*(-18) + 2y = 11
-90 + 2y = 11
2y = 101
y = 101/2